Azree an accountant, working for a foreign consultant firm and earning RM 78,000 per year is contemplating giving up his job and set up his own tax consultant firm. He estimates that renting an office would cost RM 780 per month, hiring a secretary with salary RM 1,500 per month and purchasing for required supplies would cost him RM 10,000 per annum. He estimated that his total revenues for the year would be RM 120,000. 

b)Calculate the consultant`s following cost/profit for running his own tax consultant firm:

i)         explicit cost.               

ii)        implicit cost.

iii)        accounting profit.

iv)       economic profit.      

c)Explain briefly whether Azree should practice his own tax consultant firm or not.


1
Expert's answer
2021-04-29T07:36:22-0400
"solution"

[B]calculating:

i] Explicit costs{Yearly}

"rent=RM \ 780\times 12=RM\ 9360\\
sectretary \ salary=RM\ 1,500\times12=RM\ 18000\\
supplies=RM \ 10,000"


"total \ explicit\ cost=RM \ 9360+18000+10000\\
total \ explicit\ cost =RM\ 27,360"

ii] implicit cost :

"Wages \ foregone\ /Azrees \ salary=RM\ 78,000"


iii]Accounting Profit :

"Estimated\ total\ yearly\ revenue-Total\ economic \ yearly \ profit=RM 120,000-105,360=RM14,640"


iv] Economic Profit:

"total\ implicit+explicit\ cost=RM\ 27,360+78000=RM\ 105,360"

[C]

Azree can practice his own tax consultant firm since there is economic sense in doing it as shown by the accounting profit as he could make more money/ business can be profitable.